OpenStudy (anonymous):

A person travels by car from one city to another. She drives for 35.6 min at 64.5 km/h,14.4 min at 92.3 km/h, 49.5 min at 32.4km/h,and spends 12 min along the way eating lunch and buying gas. Determine the distance between the cities along this route this is the second part of the equation i got the first part but i don't know how to do this: Determine the average speed for the trip. Answer in units of km/h

OpenStudy (kropot72):

Let d = distance between the cities in km. Let t = total time for the trip. t = 35.6 + 14.4 + 49.5 + 12.0 = 111.5 minutes = 111.5/60 hours. Then the average speed for the trip is given by: \[\large \frac{d}{\frac{111.5}{60}}=\frac{d \times60}{111.5}=you\ can\ calculate\]